OpenStudy (anonymous):

what is the radius of a sphere, circumference of a great circle, and the volume of a sphere? if the surface area of the sphere is 2304pi cm^2.. leave answers in term of pi

OpenStudy (anonymous):

radius sphere = radius of a great circle A = 4pi(r^2) 2304picm^2 = 4(r^2) 2304/4 = r^2 √576 = r radius = 24 V= (4/3)pi.r^3 = (4/3) pi .24^3 = 18.432pi

OpenStudy (anonymous):

circumference of a great circle is: 2r. pi = 2(24). pi C = 28pi